OpenStudy (anonymous):

derivative of y=3x^2sinxtanx

OpenStudy (anonymous):

\[3(2xsinxtanx+x^2cosxtanx+x^2sinxsec^2x)\] I get this using y=a'bc+ab'c+abc'

OpenStudy (asnaseer):

your answer is correct but it can be simplified a little

OpenStudy (asnaseer):

e.g.:\[\cos(x)\tan(x)=\cos(x)\times \frac{\sin(x)}{\cos(x)}=\sin(x)\]